# The foot, the fan, and the cuprate phase diagram: Fermi-volume-changing quantum phase transitions

### v5

@@ 1 Jan Zaanen and the cuprates
    A 1+1 dimensional conformal structure of the Green’s functions [16, 17]. This implied that a suitably defined relaxation time $\tau(\omega)$ obeyed Planckian scaling as a function of frequency, $\omega$, and temperature, $T$:
@@
The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [20]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the mapping of theories of quantum phase transitions in disordered metals to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 26, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], which we will discuss in Section 4. With the benefit of hindsight, we can now see some of this progress could have been made after the early papers [15, 19, 16, 18] on what was then sometimes called the SY model. But the intervening developments on the connections between SYK, black holes, and holography were important in pointing to the correct direction, and Jan played a significant role in these. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 2 The ‘foot’ and the ‘fan’
Before diving into the details of the cuprate phase diagram, it is helpful consider simpler analogous phenomena which occur in the triangular lattice antiferromagnet KYbSe2 [61]. This compound has long-range magnetic order at low $T$, and associated spin waves are observed at low energies. However, at higher energies, the neutron scattering observations (see Fig. 5) “identify a diffuse continuum with a sharp lower bound within the measured spectra …The key features of the data are reproduced by Schwinger boson theory” [61] of fractionalized spinons.
@@ 3 The ‘foot’: QPT with symmetry breaking
where $\ell=1\dots 4$ labels the 4 ordering wavevectors ${\bm{K}}_{\ell}$ at $(\pi(1\pm\delta),\pi)$ and $(\pi,\pi(1\pm\delta))$. We are interested in fluctuations of the SDW order parameters $\phi_{\ell a}$ coupled to electrons $c_{{\bm{k}}\sigma}$ with dispersion $\varepsilon({\bm{k}})$ which has a Fermi surface. So we can write down the imaginary time ($\tau$) Lagrangian
$\mathcal{L}_{1}$ $=\sum_{{\bm{k}}}c_{{\bm{k}}\sigma}^{\dagger}\left(\frac{\partial}{\partial\tau}+\varepsilon({\bm{k}})\right)c_{{\bm{k}}\sigma}+\int d^{2}{\bm{r}}\,\Bigl{\{}[s+{\color[rgb]{.75,0,.25}\delta s({\bm{r}})}][\phi({\bm{r}})]^{2}$
$+g\sum_{\ell}c^{\dagger}_{\sigma}({\bm{r}})\tau^{a}_{\sigma\sigma^{\prime}}c_{\sigma^{\prime}}({\bm{r}})\,\phi_{\ell a}({\bm{r}})e^{i{\bm{K}}_{\ell}\cdot{\bm{r}}}+K\,[\nabla_{\bm{r}}\phi({\bm{r}})]^{2}+u\,[\phi({\bm{r}})]^{4}$
@@
Here $\tau^{a}$ are the Pauli matrices, $s$ is the parameter employed to tune across the transition, and $g$ is the Yukawa coupling between the fermions and bosons. We have included two sources of spatial randomness (symbols in purple are fixed random functions of space, with no dynamics). The spatially random potential $v({\bm{r}})$, with ensemble averages $\overline{v({\bm{r}})}=0$, $\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$, acts on the fermion density, and plays a central role in the theory of disorder-induced electron localization [80]. Such fermion localization effects are also present here, but all indications are that such effects are not important for the cuprates. Instead, our focus will be on the more relevant ‘Harris disorder’, induced by spatial randomness in the position of the quantum critical point. This is represented by the spatially random ‘mass’ $\delta s({\bm{r}})$ with $\overline{\delta s({\bm{r}})}=0$, $\overline{\delta s({\bm{r}})\delta s({\bm{r}^{\prime}})}=\delta s^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$ which can lead to localization of the boson $\phi$, and so must be treated non-perturbatively.
Patel et al. [54] have recently been studied a model closely related to the 2dYSYK model (12) by large scale, high precision quantum Monte Carlo simulations (see also Ref. [53]). As we will discuss above (12), the random coupling model (12) is obtained after a mapping [43] from the random ‘mass’ model (7). Here we discuss the analysis of Ref. [49] which yields similar results: it applies a Hartree-Fock approximation to the $\phi^{4}$ term in the random mass model (7), but treats disorder numerically exactly. We integrate out the fermions (assuming fermionic eigenmodes remain extended), and considering the Landau-damped Hertz theory for the boson alone, in the presence of a random mass. We discretize on a lattice of sites (labeled by $j$), and write the SDW order parameters in terms of a real $\phi$ with a single index $a=1\ldots M$ with $M=12$. In this manner, we obtain the action
@@
The results [81, 82, 83, 84, 49, 54] show a $T=0$ FL-SDW QPT at some $s=s_{c}$, accompanied by an extended gapless quantum Griffiths phase for $s>s_{c}$, with physics similar to the random Ising model in a transverse field [86]. Fermion spectral functions and transport properties have been computed, and are qualitatively similar to the ‘foot’ of the strange metal in Fig. 3, and consistent with the recent high field observations of Campbell et al. [87] relating strange metals to spin fluctuations. Tranquada et al. [88] have also connected strange metal behavior to inhomogeneous spin fluctuations. Radaelli et al. [85] have recently observed an extended regime of critical spin fluctuations in La2-xSrxCuO4 (extending earlier observations [89, 90, 91]; see Fig. 6), and their results are in general agreement with analytic continuation in frequency of the results of Refs. [49, 54]. Similar neutron scattering results appeared recently in Ref. [92].
@@
Typical results from Ref. [49] are shown in Fig. 7. At higher $e_{\alpha}$, we obtain extended bosonic eigenmodes, where disorder can be treated in a self-averaging manner after the mapping to the 2dYSYK model to be discussed near (12). However, we observe a remarkable non-monotonic behavior in the localization length $\mathcal{L}_{\alpha}$ of the eigenmodes as $e_{\alpha}$ is reduced, which requires a non-perturbative and non-self-averaging treatment of disorder in either the random mass form of (7), or the random coupling form of (12). The bosonic modes localize at an intermediate energy, but then the localization length increases logarithmically upon further lowering the energy. This logarithmic increase is characteristic of infinite-randomness fixed points [86, 83], and arises from the interaction of the incipient localized modes. It is these localized modes of SDW spin fluctuations in Fig. 7 which are responsible for the extended critical quantum Griffiths phase, and they lead to the ‘foot’ in Fig. 3, and the critical spin fluctuations of Fig. 6 [49, 54].
@@ 4 The ‘fan’: QPT without symmetry breaking
In principle, the symmetry breaking theory for the ‘foot’ in (7) can be extended to higher temperatures to also provide a theory for the ‘fan’, and this may well be the appropriate approach for many correlated electron compounds [107]. The theory of the ‘fan’ in Refs. [43, 51] using the 2dYSYK model in (12) below also applies to the extended SDW bosons in Fig. 7.
@@
Instead the observations of Robarts et al. [109] are naturally interpreted as spinons. Note the evolution of their observations with increasing doping from a sharp spin-wave spectrum in Figs. 8a,b to a broad continuum in Figs. 8c-f. This matches the trends in the evolution of the dynamic spin structure factor of the insulating square lattice antiferromagnet with increasing second neighbor exchange, as computed by Ferrari and Becca [113] using fermionic spinons, and shown in Fig. 9. Indeed, the cuprate observations are the analog of those in the triangular lattice antiferromagnet KYbSe2 [61]. As discussed in Section 2 and Fig. 5, KYbSe2 displays a continuum at high energies which has been modeled in a theory of bosonic spinons [62, 61], along with weak antiferromagnetic order at low temperatures.
@@ 4.2 FL-FL* QPT
We now address the nature of the FL-FL* QPT in Fig. 4, in the presence of spatial disorder. We will do this for the case of the Kondo lattice transition in Fig. 10 only, as the results for the single-band model of Fig. 11 differ only in their gauge structure [57, 56, 126, 127, 130], and this is unimportant at the level of the SYK-type equations in (13). The theory is similar in structure to that for the symmetry breaking case in (7), with the Higgs field $\Phi$ replacing the SDW order parameter $\phi$: we have the imaginary time Lagrangian
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+K\,|\nabla_{\bm{r}}\Phi({\bm{r}})|^{2}+u\,|\Phi({\bm{r}})|^{4}+{\color[rgb]{.75,0,.25}v({\bm{r}})}c_{\sigma}^{\dagger}({\bm{r}})c_{\sigma}({\bm{r}})\Bigr{\}}\,,$ (10)
@@
We are interested in applying $\mathcal{L}_{2}$ to the ‘fan’ at higher temperatures, and so we can focus on the case where all the fermionic and bosonic eigenmodes are extended. However, even in the extended boson regime, a perturbative averaging over the disorder $\delta s({\bm{r}})$ in (10) does not properly account for its strong effects, and is not the correct way to proceed. The key step [43] is to apply a mapping which transfers the ‘random mass’ disorder $\delta s({\bm{r}})$ to a random Yukawa coupling disorder $g^{\prime}({\bm{r}})$ by a space dependent rescaling of $\Phi$. This has the salutary effect of accounting exactly for the leading effect of the random spatial structure of the bosonic eigenmodes, and of the Harris disorder, and placing the randomness in a form amenable to methods used in the exact solution of the SYK model [43]. In this manner we obtain the 2dYSYK theory:
$\mathcal{L}_{3}$ $=\sum_{\bm{k}}c_{{\bm{k}}\sigma}^{\dagger}\left(\frac{\partial}{\partial\tau}+\varepsilon({\bm{k}})\right)c_{{\bm{k}}\sigma}+\sum_{\bm{k}}f_{{\bm{k}}\sigma}^{\dagger}\left(\frac{\partial}{\partial\tau}+\varepsilon_{1}({\bm{k}})\right)f_{{\bm{k}}\sigma}$
[…]
The coupling $g^{\prime}({\bm{r}})$ obeys $\overline{g^{\prime}({\bm{r}})}=0$, $\overline{g^{\prime}({\bm{r}})g^{\prime}({\bm{r}^{\prime}})}=g^{\prime 2}\delta({\bm{r}}-{\bm{r}^{\prime}})$. The theory $\mathcal{L}_{3}$ has been studied perturbatively in $g$ and $\color[rgb]{.75,0,.25}g^{\prime}$ in Ref. [43]. After adding a large number of flavors and making the couplings random in flavor space, as in the SYK model, we can obtain the fully self-consistent solution SYK-type equations [43]:
@@
A theory closely related to (12) has also been studied using non-pertubative quantum Monte Carlo at $g=0$ in Ref. [54]. Their results at low $T$ display a localized boson foot regime, similar to that of the random mass model (8) in Ref. [49]. This agreement is evidence for the validity of the mapping from (10) to (12).
@@ 5 Confinement crossover
where $(B_{i1},B_{i2})$ are charge $e$ spinless bosons. This description has an emergent SU(2) gauge field, and $B$ is a SU(2) fundamental Higgs field [134]. This approach has been developed for a theory of the pseudogap as a FL* with a $\pi$-flux spin liquid, and its lower $T$ confinement transitions in Refs. [131, 135, 136].
@@
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for future transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [141], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet. The matching of ADMR to the ordered antiferromagnet Fermi surface [114] therefore supports the holon metal.

### v6

@@ 1 Jan Zaanen and the cuprates
    A 1+1 dimensional conformal structure of the Green’s functions [16, 17] in a compressible system. This implied that a suitably defined relaxation time $\tau(\omega)$ obeyed Planckian scaling as a function of frequency, $\omega$, and temperature, $T$:
@@
The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [20]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the mapping of theories of quantum phase transitions in disordered metals to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 26, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], which we will discuss in Sections 3 and 4. With the benefit of hindsight, we can now see some of this progress could have been made after the early papers [15, 19, 16, 18] on what was then sometimes called the SY model. But the intervening developments on the connections between SYK, black holes, and holography were important in pointing to the correct direction, and Jan played a significant role in these. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 2 The ‘foot’ and the ‘fan’
Before diving into the details of the cuprate phase diagram, it is helpful to consider simpler analogous phenomena which occur in the triangular lattice antiferromagnet KYbSe2 [61]. This compound has long-range magnetic order at low $T$, and associated spin waves are observed at low energies. However, at higher energies, the neutron scattering observations (see Fig. 5) “identify a diffuse continuum with a sharp lower bound within the measured spectra …The key features of the data are reproduced by Schwinger boson theory” [61] of fractionalized spinons.
@@ 3 The ‘foot’: QPT with symmetry breaking
where $\ell=1\dots 4$ labels the 4 ordering wavevectors ${\bm{K}}_{\ell}$ at $(\pi(1\pm\delta),\pi)$ and $(\pi,\pi(1\pm\delta))$. We are interested in fluctuations of the SDW order parameters $\phi_{\ell a}$ coupled to electrons $c_{{\bm{k}}\sigma}$ with dispersion $\varepsilon({\bm{k}})$ which has a Fermi surface. We describe this with a 2dYSYK model [41, 43] with imaginary time ($\tau$) Lagrangian
$\mathcal{L}_{1}$ $=\sum_{{\bm{k}}}c_{{\bm{k}}\sigma}^{\dagger}\left(\frac{\partial}{\partial\tau}+\varepsilon({\bm{k}})\right)c_{{\bm{k}}\sigma}+\int d^{2}{\bm{r}}\,\Bigl{\{}s\,[\phi({\bm{r}})]^{2}$
$+\sum_{\ell}[g+{\color[rgb]{.75,0,.25}g^{\prime}({\bm{r}})}]c^{\dagger}_{\sigma}({\bm{r}})\tau^{a}_{\sigma\sigma^{\prime}}c_{\sigma^{\prime}}({\bm{r}})\,\phi_{\ell a}({\bm{r}})e^{i{\bm{K}}_{\ell}\cdot{\bm{r}}}+K\,[\nabla_{\bm{r}}\phi({\bm{r}})]^{2}+u\,[\phi({\bm{r}})]^{4}$
@@
Here $\tau^{a}$ are the Pauli matrices, $s$ is the parameter employed to tune across the transition, and $g$ is the Yukawa coupling between the fermions and bosons. We have included two sources of spatial randomness (symbols in purple are fixed random functions of space, with no dynamics). The spatially random potential $v({\bm{r}})$, with ensemble averages $\overline{v({\bm{r}})}=0$, $\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$, acts on the fermion density, and plays a central role in the theory of disorder-induced electron localization [80]. Such fermion localization effects are also present here, but all indications are that such effects are not important for the cuprates.
Instead, our focus will be on the more relevant ‘Harris disorder’, induced by spatial randomness in the position of the quantum critical point. This is represented in (7) by the spatially disordered Yukawa coupling $g^{\prime}({\bm{r}})$ with $\overline{g^{\prime}({\bm{r}})}=0$, $\overline{g^{\prime}({\bm{r}})g^{\prime}({\bm{r}^{\prime}})}=g^{\prime 2}\delta({\bm{r}}-{\bm{r}^{\prime}})$. A more conventional form of the Harris disorder is in terms of a ‘random mass’ under which the coupling $s\rightarrow s+{\color[rgb]{.75,0,.25}\delta s({\bm{r}})}$; but this can be mapped to $g^{\prime}({\bm{r}})$ by a rescaling of $\phi$ chosen to make the co-efficient of $[\phi({\bm{r}})]^{2}$ independent of ${\bm{r}}$. In either form, the Harris disorder leads to localization of the boson $\phi$ at low energies (as we discuss below), and so must be treated non-perturbatively. However, in the higher energy regime where the bosons are extended, it is preferable to place the Harris disorder only in $g^{\prime}({\bm{r}})$ as this has the salutary effect of accounting for the spatial structure of the bosonic eigenmodes, and so enables use of the self-consistent SYK methods for the spatially averaged Green’s functions.
Patel et al. [54] have recently studied the 2dYSYK model (7) at $g=0$ by large scale, high precision quantum Monte Carlo simulations (see also Ref. [53]). Here we discuss the analysis of Ref. [49] which yields similar results: it applies a Hartree-Fock approximation to the $\phi^{4}$ term in an Hertz effective theory for the bosons alone, but treats disorder numerically exactly. We integrate out the fermions from (7) (assuming fermionic eigenmodes remain extended), and consider the Landau-damped Hertz theory for the boson alone. The disorder in $g^{\prime}({\bm{r}})$ will lead to spatial disorder in all couplings in the Hertz theory, but we retain only the most relevant ‘random mass’ disorder $\delta s({\bm{r}})$. We discretize on a lattice of sites (labeled by $j$), and write the SDW order parameters in terms of a real $\phi$ with a single index $a=1\ldots M$ with $M=12$. In this manner, we obtain the action
@@
The results [81, 82, 83, 84, 49, 54] show a $T=0$ FL-SDW QPT at some $s=s_{c}$, accompanied by an extended gapless quantum Griffiths phase for $s>s_{c}$, with physics similar to the random Ising model in a transverse field [86]. Fermion spectral functions and transport properties have been computed, and are qualitatively similar to the ‘foot’ of the strange metal in Fig. 3. We note that a similar localized boson foot regime has also been found in the non-pertubative quantum Monte Carlo at $g=0$ [54].
These results are consistent with the recent high field observations of Campbell et al. [87] relating strange metals to spin fluctuations. Tranquada et al. [88] have also connected strange metal behavior to inhomogeneous spin fluctuations. Radaelli et al. [85] have recently observed an extended regime of critical spin fluctuations in La2-xSrxCuO4 (extending earlier observations [89, 90, 91]; see Fig. 6), and their results are in general agreement with analytic continuation in frequency of the results of Refs. [49, 54]. Similar neutron scattering results appeared recently in Ref. [92].
@@
Typical results from Ref. [49] are shown in Fig. 7. At higher $e_{\alpha}$, we obtain extended bosonic eigenmodes, where disorder can be treated in a self-averaging manner, as will be discussed near (12). However, we observe a remarkable non-monotonic behavior in the localization length $\mathcal{L}_{\alpha}$ of the eigenmodes as $e_{\alpha}$ is reduced, which requires a non-perturbative and non-self-averaging treatment of disorder in either the random mass form of (7), or the random coupling form of (10). The bosonic modes localize at an intermediate energy, but then the localization length increases logarithmically upon further lowering the energy. This logarithmic increase is characteristic of infinite-randomness fixed points [86, 83], and arises from the interaction of the incipient localized modes. It is these localized modes of SDW spin fluctuations in Fig. 7 which are responsible for the extended critical quantum Griffiths phase, and they lead to the ‘foot’ in Fig. 3, and the critical spin fluctuations of Fig. 6 [49, 54].
@@ 4 The ‘fan’: QPT without symmetry breaking
In principle, the symmetry breaking theory for the ‘foot’ in (7) can be extended to higher temperatures to also provide a theory for the ‘fan’, and this may well be the appropriate approach for many correlated electron compounds [107]. The theory of the ‘fan’ in Refs. [43, 51] using the disorder averaged in (12) below also applies to the extended SDW bosons in Fig. 7.
@@
Instead the observations of Robarts et al. [109] are naturally interpreted as spinons. Note the evolution of their observations with increasing doping from a sharp spin-wave spectrum in Figs. 8a,b to a broad continuum in Figs. 8c-f. This matches Fig. 9, which shows the evolution of the dynamic spin structure factor of the insulating square lattice antiferromagnet with increasing second neighbor exchange $J_{2}$, as computed by Ferrari and Becca [113] using fermionic spinons (see also the connection of spinons in the unfrustrated square lattice to the quasi-one-dimensional limit [114, 115] . There is also a correspondence in the energy scale, measured in units of the first neighbor exchange $J_{1}\approx 130$ meV [116]. Indeed, the cuprate observations are the analog of those in the triangular lattice antiferromagnet KYbSe2 [61]. As discussed in Section 2 and Fig. 5, KYbSe2 displays a continuum at high energies which has been modeled in a theory of bosonic spinons [62, 61], along with weak antiferromagnetic order at low temperatures.
@@ 4.2 FL-FL* QPT
We now address the nature of the FL-FL* QPT in Fig. 4, in the presence of spatial disorder. We will do this for the case of the Kondo lattice transition in Fig. 10 only, as the results for the single-band model of Fig. 11 differ only in their gauge structure [57, 56, 129, 130, 133], and this is unimportant at the level of the SYK-type equations in (12). The theory is similar in structure to that for the symmetry breaking case in (7), with the Higgs field $\Phi$ replacing the SDW order parameter $\phi$: we have the imaginary time Lagrangian for a 2dYSYK theory:
@@
We are interested in applying $\mathcal{L}_{2}$ to the ‘fan’ at higher temperatures, and so we can focus on the case where all the fermionic and bosonic eigenmodes are extended, and we apply self-averaging SYK techniques. After adding a large number of flavors and making the couplings random in flavor space, as in the SYK model, we can obtain the fully self-consistent solution SYK-type equations [41, 43]:
@@ 5 Confinement crossover
where $(B_{i1},B_{i2})$ are charge $e$ spinless bosons. This description has an emergent SU(2) gauge field, and $B$ is an SU(2) fundamental Higgs field [137]. This approach has been developed for a theory of the pseudogap as a FL* with a $\pi$-flux spin liquid, and its lower $T$ confinement transitions in Refs. [134, 138, 139].
@@
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for future transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [144], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet.